# MathMathgg

### v1

@@ Abstract
We present an ab initio calculation of nuclear $\beta$-decay within the framework of nuclear lattice effective field theory (NLEFT), employing auxiliary-field quantum Monte Carlo methods to solve the nuclear many-body problem. Our approach combines next-to-next-to-leading order (NNLO) two- and three-body chiral interactions with one- and two-body axial current operators, all consistently derived in chiral effective field theory. Low-energy constants are determined exclusively from nucleon-nucleon scattering phase shifts and few-body observables for systems with $A\leq 3$. Using these interactions and transition operators, we perform two-channel Monte Carlo simulations to compute the $\beta$-decay matrix element for 6He, obtaining results in reasonable agreement with experimental measurements. To address the Monte Carlo sign problem, we implement a perturbative expansion around a leading-order Hamiltonian with approximate Wigner-SU(4) symmetry. This systematic approach provides a foundation for extending NLEFT simulations to precision studies of weak processes in medium-mass nuclei.
@@ I Introduction
Nuclear $\beta$-decay is a fundamental process crucial to understanding the laws of nature. Historically, studies of $\beta$-decay have led to groundbreaking discoveries, including the existence of neutrinos Pauli (1978), the confirmation of parity violation Wu et al. (1957), and contributions to the unification of electroweak interactions Glashow (1961); Weinberg (1967); Salam (1968). Today, nuclear $\beta$-decay provides a sensitive probe for physics beyond the Standard Model. For a comprehensive review of ongoing and planned experimental programs, see Ref. González-Alonso et al. (2019) and references therein.
@@
Besides its central role in particle physics, nuclear $\beta$-decay is also of great interest within the field of low-energy nuclear physics. Currently there remain several unresolved challenges in precisely describing the nuclear weak decay with the nucleonic degrees of freedom. For instance, traditional shell-model calculations tend to overestimate the reduced Gamow-Teller matrix element (RGTME), especially for medium-mass nuclei Brown and Wildenthal (1985); Martínez-Pinedo et al. (1996); Langanke et al. (1995). To reconcile this discrepancy, the axial charge $g_{A}$ in the axial current is often tuned to match experimental data Chou et al. (1993), leading to the well-known $g_{A}$ quenching problem. Moreover, the nuclear force binds nucleons into atomic nuclei, induces complicated many-body correlations, which requires tremendous amount of computing resources and advanced algorithms to solve the corresponding quantum many-body problems. Lastly, the external electroweak field interacts with the nucleus via not only one-body currents but also many-body currents, so a proper description of $\beta$-decays requires a consistent treatment of the electroweak currents together with the complicated nuclear chiral interactions.
@@
Recent years have seen rapid development of $ab$ $initio$ nuclear many-body methods Carlson et al. (2015); Barrett et al. (2013); Hagen et al. (2014); Hergert et al. (2016); Lee (2009), which provide a consistent and systematically improvable framework for understanding the intricate interplays among nuclear forces, electroweak currents and quantum many-body correlations in $\beta$-decay processes. For nuclei of $s$ and $p$ shell, calculations based on the quantum Monte Carlo (QMC) method Pastore et al. (2018); King et al. (2020) and the no core shell model method Navrátil et al. (2007); Gysbers et al. (2019) have been performed systematically, with the latter applied to the study of the anomalous long lifetime of 14Maris et al. (2011). For heavier nuclei, the coupled-cluster method has been applied to investigating the impact of three-body forces and two-body currents on the Gamow-Teller transitions of 14C, 22O and 24Ekström et al. (2014), and it also provided the first $ab$ $inito$ calculation of 100Sn $\beta$-decay Gysbers et al. (2019). Besides, the valence-space in-medium similarity renormalization group method has made significant progress in studying $\beta$-decays of medium-mass nuclei Gysbers et al. (2019); Stroberg (2021); Yuan et al. (2025). These $ab$ $initio$ calculations have significantly improved our understanding of nuclear $\beta$-decays, yet also raised a lot of open questions. For instance, the QMC calculations struggle to reproduce the experimental $\beta$-decay rate for 8Li, 8B, and 8He King et al. (2020). Furthermore, the contribution of two-body currents is found to have opposite signs in Refs. King et al. (2020) and Gysbers et al. (2019). This theoretical inconsistency was further explored in Ref. Gnech et al. (2021), which also revealed that the one-pion-exchange (OPE) current is highly sensitive to the nuclear forces and regularization schemes.
@@
As one of the advanced nuclear ab initio methods, the nuclear lattice effective field theory (NLEFT) Lee (2009); Lähde and Meißner (2019) implements the chiral interactions on a cubic spatial lattice and solves the corresponding discrete many-body Schrödinger equation using auxiliary-field Monte Carlo method. Recently the NLEFT has achieved remarkable progress in studying the properties of nuclear ground states and excited states Borasoy et al. (2007); Epelbaum et al. (2010a, b); Lähde et al. (2014); Lu et al. (2019, 2022); Elhatisari et al. (2024a); Epelbaum et al. (2014); Shen et al. (2023); Meißner et al. (2024); Shen et al. ; Zhang et al. ; Wang et al. , nuclear clustering Epelbaum et al. (2011, 2012, 2013); Elhatisari et al. (2016, 2017), scattering processes Bour et al. (2012); Elhatisari et al. (2015), thermal nuclear matter Lu et al. (2020); Ren et al. (2024); Ma et al. (2024a) , and hypernucleus Bour et al. (2015); Frame et al. (2020); Hildenbrand et al. (2024). In this area, recent years a significant amount of efforts was made to build high-precision lattice interactions and design algorithms that mitigate the Monte Carlo sign problem. For example, the wave function matching method Elhatisari et al. (2024a) and perturbative quantum Monte Carlo method Lu et al. (2022); Liu et al. were proposed to improve the accuracy of perturbative calculations around simple Hamiltonian without the sign problem. When combined with high-fidelity nuclear forces, these methods were able to reproduce the experimental binding energies and charge radii with systematically controllable statistical errors.
Attempts to study nuclear $\beta$-decay through NLEFT have just begun very recently. The lifetime of triton $\beta$-decay is calculated in Ref. Elhatisari et al. (2024b) using the Lanczos eigenvector method Lanczos (1950) to diagonalize the lattice Hamiltonian directly. However, for nuclei with mass number $A\geq 4$ the direct diagonalization is not possible due to the exponential increase of the computational resources. In this case we need the quantum Monte Carlo techniques which have polynomial scaling if the sign problem is well under control. Furthermore, the currents beyond leading order (LO) are omitted in Ref. Elhatisari et al. (2024b), leaving their effects undetected. In this regard, it is desirable to improve and apply the NLEFT Monte Carlo approach to investigating $\beta$-decays of light and medium-mass nuclei. The results would also provide an alternative benchmark for calculations with other ab initio methods.
In this work, we address this demand by presenting an in-depth study of nuclear $\beta$-decay within the NLEFT framework, focusing on the processes $\rightarrow{{}^{3}}$e and 6He $\rightarrow{{}^{6}}$i. The former serves as a benchmark for validating our methodology, while the latter illustrates its potential applicability to $\beta$-decays beyond few-body sector. We use next-to-next-to-leading order (NNLO) chiral interactions alongside nuclear axial currents up to next-to-next-to-next-to-leading order (N3LO), consistently derived from chiral EFT, to compute the RGTME. Since this is an exploratory study, our primary goal is to develop the necessary computational techniques and assess the reasonableness of our predictions. A more systematic study of $\beta$-decays in even heavier nuclei is left for the future.
@@
This paper is organized as follows. The nuclear interaction and axial currents used in this work are described in Sec. II and III. Our calculation methods and techniques are briefly discussed in Sec. IV. Results for 3H and 6He $\beta$-decay are presented in Sec. V. More details are provided in the appendix.
@@ II.1 Full Interaction
The full Hamiltonian used in this work is the NNLO chiral Hamiltonian discretized on a lattice with a lattice spacing $a=1.64$ fm,
@@
The numerical values of the parameters used are $\tilde{g}_{A}=1.287$ for the nucleon axial charge, $f_{\pi}=92.2$ MeV for the pion decay constant, $M_{\pi}=134.98$ MeV for the pion mass, and $\Lambda_{\pi}=300$ MeV for the local regulator. The axial charge $\tilde{g}_{A}$ incorporates the Goldberger-Treiman discrepancy Arndt et al. (1994) and differs slightly from value $g_{A}=1.2723(23)$ Tanabashi et al. (2018) in the axial current. To distinguish between them, we use a tilde notation. The term in Eq. (3) serves as a counterterm to remove the short-distance singularity in the OPE potential. It is worth noting that the characteristic momentum scale in this work is around 100 MeV, considerably lower than twice the pion mass. As a result, the two-pion exchange (TPE) potential, which first appears at NLO, cannot be distinguished from contact terms Epelbaum et al. (2009).
@@
where $B_{i}$ and $C_{i}$ are low-energy constants (LECs) fitted to the Nijmegen neutron-proton ($n$-$p$) scattering phase shifts Stoks et al. (1993). All contact terms are regularized using a nonlocal Gaussian regulator in momentum space, following Ref. Lu and Deng . Specifically, for the matrix element $V(\bm{p}_{1},\bm{p}_{1}^{\prime},\bm{p}_{2},\bm{p}_{2}^{\prime})$, where $\bm{p}_{i}$ and $\bm{p}^{\prime}_{i}$ are the incoming and outgoing momenta of the $i$-th nucleon, we apply a non-local momentum regulator
@@
These isospin-breaking terms are essential for reproducing the 3H-3He mass splitting Epelbaum et al. (2010b). The 2N LECs $C_{pp}$ and $C_{nn}$ in Eqs. (8)-(9) are fitted to experimental $p$-$p$ phase shifts, as well as $n$-$n$ scattering lengths and effective ranges Gonzalez Trotter et al. (2006); Miller et al. (1990). Details on determining the LECs are provided in Appendix A.
Finally, for the three-nucleon (3N) potential $V_{3\mathrm{N}}^{Q^{3}}$, which first appears at NNLO, we include the OPE term
@@
II.2 Non-perturbative Interaction
The full Hamiltonian in Eq. (1) includes complex spin-isospin-dependent terms, which pose significant challenges for lattice calculations. One major issue is that these terms can induce severe sign problems in fermionic systems, such as nucleons, in Monte Carlo simulations Lähde and Meißner (2019). To mitigate this, one can construct a simplified Hamiltonian $H_{0}$ which captures the essential physics but with much milder sign problems, and solve the full Hamiltonian $H$ perturbatively Lu et al. (2022). Details of the perturbative method will be provided in Sec. IV.1. In this work, we consider the following non-perturbative Hamiltonian:
@@ II.2 Non-perturbative Interaction
where $s$ and $i$ are the spin and isospin index, respectively. The parameter $s_{\mathrm{L}}$ controls the range of the local part of the interaction. The smeared creation operator is defined as
@@
with $\tau_{I}$ the $I$-th Pauli matrix .
The Hamiltonian in Eq. (13) consists of two parts, written separately for clarity. The first line contains terms that respect Wigner’s SU(4) symmetry Wigner (1937), an approximate symmetry of low-energy nuclear interactions. The inclusion of this symmetry helps suppress sign problems in lattice Monte Carlo calculations Chen et al. (2004). Recently, the SU(4)-symmetric Hamiltonian has been applied to the study of the properties of light and medium-mass nuclei as well as the equation of state of symmetric nuclear matter and neutron matter Lu et al. (2019), the intrinsic clustering geometry of 12C’s low-lying states Shen et al. (2023), the thermodynamics of nuclear systems Lu et al. (2020), the clustering effects in hot, dilute nuclear matter Ren et al. (2024), and the alpha-particle monopole transition form factor Meißner et al. (2024). The second line of Eq. (13) introduces an isospin-dependent contact term and a OPE term, both of which explicitly break the SU(4) symmetry. These additional terms are introduced to improve the overall description of $p$-shell nuclei Wiringa and Pieper (2002) , such as 6He and 6Li studied in this work. To suppress the sign problem caused by the tensor force, we adopt a lower cutoff of $\Lambda_{\pi}^{\prime}=180$ MeV for the OPE term in Eq. (13), compared to the higher cutoff of $\Lambda_{\pi}=300$ MeV used in the full Hamiltonian in Eq. (1).
@@ III Axial Current
Figure 1: Diagrams illustrating the contributions to the axial current up to N3LO used in this work. Nucleons, pions and the external field are denoted by the solid , dashed, and wavy lines respectively. Panel (a) denotes the LO current while Panel (b) is the relativistic correction at NNLO. Panel (c) and Panel (d) are the OPE current and the contact current at N3LO, respectively.
@@
Since the momentum carried by the current is very small in $\beta$-decay processes, we work in the limit of vanishing current momenta to simplify calculations. The explicit momentum-space (MS) expressions of the axial currents under this limit are provided in Ref. Baroni et al. (2016b), and a diagrammatic representation of the currents is shown in Fig. 1. For two-body currents at N3LO, we regularize them using the regulator $f_{2\mathrm{N}}$ introduced in Eq. (5). The N3LO OPE current involves the two $N$-$\pi$ coupling constants, $c_{3}$ and $c_{4}$, for which we adopt the values $c_{3}=-5.61$ GeV-1 and $c_{4}=4.26$ GeV-1 from Ref. Hoferichter et al. (2015).
Here, we highlight a subtle difference between our calculation on a discretized cubic lattice and previous studies in the continuum Park et al. (2003); Baroni et al. (2016b, 2018); Gnech et al. (2021); King et al. (2020). In the latter case, the MS expression of the current is first transformed into configuration space (CS) analytically, i.e., $\bm{J}_{5}^{\mathrm{MS}}\rightarrow\bm{J}_{5}^{\mathrm{CS}}$. In this step, some parts of the OPE current, $\bm{J}_{5}^{\mathrm{MS}}(\mathrm{OPE})$, are absorbed into the definition of the contact current, $\bm{J}_{5}^{\mathrm{CS}}(\mathrm{CT})$, using Fermi-Dirac statistics Park et al. (2003). As a result, if the MS expression of the contact term is
$\bm{J}^{\mathrm{MS}}_{5,a}(\mathrm{CT})=z_{0}(\bm{\tau}_{1}\times\bm{\tau}_{2})_{a}\bm{\sigma}_{1}\times\bm{\sigma}_{2},$ (18)
then its CS expression takes the form
$\bm{J}^{\mathrm{CS}}_{5,a}(\mathrm{CT})=z^{\prime}_{0}(\bm{\tau}_{1}\times\bm{\tau}_{2})_{a}\bm{\sigma}_{1}\times\bm{\sigma}_{2}.$ (19)
The two coefficients $z_{0}$ and $z_{0}^{\prime}$ are related to each other through
$z_{0}^{\prime}\propto z_{0}+\frac{1}{6}c_{3}+\frac{1}{3}c_{4}+\frac{1}{12M_{N}},$ (20)
with the last three terms $c_{3}/6$, $c_{4}/3$, and $1/(12M_{N})$ originally belonging to $\bm{J}_{5}^{\mathrm{MS}}(\mathrm{OPE})$. However, in our lattice calculation, since both the momenta and the coordinates are discretized, the transformation from MS to CS can be efficiently handled numerically using fast Fourier transform, and these is no need to perform any analytical calculation. Therefore, we do not split the OPE current into two parts, and the MS and CS currents remain in one-to-one correspondence, eliminating any ambiguity. Finally, we note that the coefficient $z_{0}$ in Eq. (18) is related to the three-body LEC $c_{D}$ via Baroni et al. (2018)
$z_{0}=-\frac{c_{D}}{8\Lambda_{\chi}f_{\pi}^{2}},$ (21)
and we take $\Lambda_{\chi}=700$ MeV for the consistency with Eq (11).

### v2

# Investigating nuclear $\beta$ decay using a lattice quantum Monte Carlo approach

@@ Abstract
We present an ab initio calculation of nuclear $\beta$ decay within the framework of nuclear lattice effective field theory (NLEFT), employing auxiliary-field quantum Monte Carlo methods to solve the nuclear many-body problem. Our approach combines next-to-next-to-leading order two- and three-body chiral interactions with one- and two-body axial current operators, all consistently derived in chiral effective field theory. Low-energy constants are determined exclusively from nucleon-nucleon scattering phase shifts and few-body observables for systems with $A\leq 3$. Using these interactions and transition operators, we perform two-channel Monte Carlo simulations to compute the $\beta$-decay matrix element for 6He, obtaining results in reasonable agreement with experimental measurements. To address the Monte Carlo sign problem, we implement a perturbative expansion around a leading-order Hamiltonian with approximate Wigner-SU(4) symmetry. This systematic approach provides a foundation for extending NLEFT simulations to precision studies of weak processes in medium-mass nuclei.
@@ I Introduction
Nuclear $\beta$ decay is a fundamental process crucial to understanding the laws of nature. Historically, studies of $\beta$ decay have led to groundbreaking discoveries, including the existence of neutrinos Brown (1978), the confirmation of parity violation Wu et al. (1957), and contributions to the unification of electroweak interactions Glashow (1961); Weinberg (1967); Salam (1968). Today, nuclear $\beta$ decay provides a sensitive probe for physics beyond the standard model. For a comprehensive review of ongoing and planned experimental programs, see Ref. González-Alonso et al. (2019) and references therein.
@@
Besides its central role in particle physics, nuclear $\beta$ decay is also of great interest within the field of low-energy nuclear physics. Currently there remain several unresolved challenges in precisely describing the nuclear weak decay with the nucleonic degrees of freedom. For instance, traditional shell-model calculations tend to overestimate the reduced Gamow-Teller matrix element (RGTME), especially for medium-mass nuclei Brown and Wildenthal (1985); Martínez-Pinedo et al. (1996); Langanke et al. (1995). To reconcile this discrepancy, the axial charge $g_{A}$ in the axial current is often tuned to match experimental data Chou et al. (1993), leading to the well-known $g_{A}$ quenching problem. Moreover, the nuclear force binds nucleons into atomic nuclei, inducing complicated many-body correlations, which requires tremendous amount of computing resources and advanced algorithms to solve the corresponding quantum many-body problems. Lastly, the external electroweak field interacts with the nucleus via not only one-body currents but also many-body currents, so a proper description of $\beta$ decays requires a consistent treatment of the electroweak currents together with the complicated nuclear chiral interactions.
@@
Recent years have seen rapid development of abinitioab\ initio nuclear many-body methods Carlson et al. (2015); Barrett et al. (2013); Hagen et al. (2014); Hergert et al. (2016); Lee (2009), which provide a consistent and systematically improvable framework for understanding the intricate interplays among nuclear forces, electroweak currents, and quantum many-body correlations in $\beta$-decay processes. For nuclei of $s$ and $p$ shells, calculations based on the quantum Monte Carlo (QMC) method Pastore et al. (2018); King et al. (2020) and the no-core shell model method Navrátil et al. (2007); Gysbers et al. (2019) have been performed systematically, with the latter applied to the study of the anomalous long lifetime of 14Maris et al. (2011). For heavier nuclei, the coupled-cluster method has been applied to investigating the impact of three-body forces and two-body currents on the Gamow-Teller transitions of 14C, 22O and 24Ekström et al. (2014), and it also provided the first abab initoinito calculation of 100Sn $\beta$ decay Gysbers et al. (2019). Besides, the valence-space in-medium similarity renormalization group method has made significant progress in studying $\beta$ decays of medium-mass nuclei Gysbers et al. (2019); Stroberg (2021); Yuan et al. (2025). These abab initioinitio calculations have significantly improved our understanding of nuclear $\beta$ decays, yet also raised a lot of open questions. For instance, the QMC calculations struggle to reproduce the experimental $\beta$ decay rate for 8Li, 8B, and 8He King et al. (2020). Furthermore, the contribution of two-body currents is found to have opposite signs in Refs. King et al. (2020) and Gysbers et al. (2019). This theoretical inconsistency was further explored in Ref. Gnech et al. (2021), which also revealed that the one-pion-exchange (OPE) current is highly sensitive to the nuclear forces and regularization schemes.
@@
As one of the advanced nuclear ab initio methods, the nuclear lattice effective field theory (NLEFT) Lee (2009); Lähde and Meißner (2019) implements the chiral interactions on a cubic spatial lattice and solves the corresponding discrete many-body Schrödinger equation using the auxiliary-field Monte Carlo method. Recently the NLEFT has achieved remarkable progress in studying the properties of nuclear ground states and excited states Borasoy et al. (2007); Epelbaum et al. (2010a, b); Lähde et al. (2014); Lu et al. (2019, 2022); Elhatisari et al. (2024a); Epelbaum et al. (2014); Shen et al. (2023); Meißner et al. (2024); Shen et al. (2025); Zhang et al. ; Wang et al. , nuclear clustering Epelbaum et al. (2011, 2012, 2013); Elhatisari et al. (2016a, 2017), scattering processes Bour et al. (2012); Elhatisari et al. (2015), thermal nuclear matter Lu et al. (2020); Ren et al. (2024); Ma et al. (2024a) , and hypernuclei Bour et al. (2015); Frame et al. (2020); Hildenbrand et al. (2024). In this area, in recent years a significant amount of effort was made to build high-precision lattice interactions and design algorithms that mitigate the Monte Carlo sign problem. For example, the wave function matching method Elhatisari et al. (2024a) and perturbative quantum Monte Carlo method Lu et al. (2022); Liu et al. (2025) were proposed to improve the accuracy of perturbative calculations around a simple Hamiltonian without the sign problem. When combined with high-fidelity nuclear forces, these methods were able to reproduce the experimental binding energies and charge radii with systematically controllable statistical errors.
Attempts to study nuclear $\beta$ decay through NLEFT have just begun very recently. The first and currently the only study focuses on the triton $\beta$ decay Elhatisari et al. (2024b), which employs the recently developed high-fidelity next-to-next-to-next-to-leading order (N3LO) chiral interaction Elhatisari et al. (2024a) and the leading-order (LO) axial current to achieve consistent predictions for the Fermi and Gamow-Teller matrix elements. The sensitivity of these matrix elements to the strengths of three-body forces is also carefully examined therein.
In this work, we extend the work of Elhatisari et al. (2024b) by presenting an in-depth NLEFT study of nuclear $\beta$ decay, with major developments of two aspects. First, we have incorporated one- and two-body axial currents beyond LO in the calculation, which are essential for the determination of three-body low-energy constants (LECs), $c_{D}$ and $c_{E}$, of the next-to-next-to-leading order (N2LO) chiral nuclear force through triton $\beta$ decay  Gazit et al. (2009), and it has been shown that the two-body currents provide sizable contributions to Gamow-Teller matrix elements of certain nuclei King et al. (2020); Gysbers et al. (2019). Second, we have succeeded in simulating nuclear $\beta$ decay through the lattice quantum Monte Carlo method, which can be directly applied to weak decays of medium-mass and heavy nuclei. As a first step, here we focus on the two simplest $\beta$ decay processes, i.e., 3H $\rightarrow{{}^{3}}$He and 6He $\rightarrow{{}^{6}}$Li. We first fix $c_{D}$ and $c_{E}$ through the former process, then validate the applicability of our method using the latter. Since the main goal of this work is to build the relevant Monte Carlo machinery for investigating nuclear $\beta$ decay through NLEFT, we use the N2LO lattice chiral interaction fitted to nucleon-nucleon scattering phase shifts and few-body observables with $A\leq 3$ for numerical convenience. A systematic study based on the high-fidelity N3LO interaction is left for the future.
@@
This paper is organized as follows. The nuclear interaction and axial currents used in this work are described in Secs. II and III. Our calculation methods and techniques are briefly discussed in Sec. IV. Results for 3H and 6He $\beta$ decay are presented in Sec. V. More details are provided in the Appendixes.
@@ II.1 Full interaction
The full Hamiltonian used in this work is the N2LO chiral Hamiltonian discretized on a lattice with a lattice spacing $a=1.64$ fm,
@@
The numerical values of the parameters used are $\tilde{g}_{A}=1.287$ for the nucleon axial charge, $f_{\pi}=92.2$ MeV for the pion decay constant, $M_{\pi}=134.98$ MeV for the pion mass, and $\Lambda_{\pi}=300$ MeV for the local regulator. The axial charge $\tilde{g}_{A}$ incorporates the Goldberger-Treiman discrepancy Arndt et al. (1994) and differs slightly from value $g_{A}=1.2723(23)$ Tanabashi et al. (2018) in the axial current. To distinguish between them, we use a tilde notation. The term in Eq. (3) serves as a counterterm to remove the short-distance singularity in the OPE potential. It is worth noting that the characteristic momentum scale in this work is around 100 MeV, considerably lower than twice the pion mass. As a result, the two-pion exchange (TPE) potential, which first appears at next-to leading order (NLO), cannot be distinguished from contact terms Epelbaum et al. (2009).
@@
where $B_{i}$ and $C_{i}$ are LECs fitted to the Nijmegen neutron-proton ($n$-$p$) scattering phase shifts Stoks et al. (1993). All contact terms are regularized using a nonlocal Gaussian regulator in momentum space, following Ref. Lu and Deng . Specifically, for the matrix element $V(\bm{p}_{1},\bm{p}_{1}^{\prime},\bm{p}_{2},\bm{p}_{2}^{\prime})$, where $\bm{p}_{i}$ and $\bm{p}^{\prime}_{i}$ are the incoming and outgoing momenta of the $i$th nucleon, we apply a non-local momentum regulator
@@
These isospin-breaking terms are essential for reproducing the 3H-3He mass splitting Epelbaum et al. (2010b). The 2N LECs $C_{pp}$ and $C_{nn}$ in Eqs. (8) and (9) are fitted to experimental $p$-$p$ phase shifts, as well as $n$-$n$ scattering lengths and effective ranges Gonzalez Trotter et al. (2006); Miller et al. (1990). Details on determining the LECs are provided in Appendix A.
Finally, for the three-nucleon (3N) potential $V_{3\mathrm{N}}^{Q^{3}}$, which first appears at N2LO, we include the OPE term
@@
II.2 Nonperturbative interaction
The full Hamiltonian in Eq. (1) includes complex spin-isospin-dependent terms, which pose significant challenges for lattice calculations. One major issue is that these terms can induce severe sign problems in fermionic systems, such as nucleons, in Monte Carlo simulations Lähde and Meißner (2019). To mitigate this, one can construct a simplified Hamiltonian $H_{0}$ which captures the essential physics but with much milder sign problems, and solve the full Hamiltonian $H$ perturbatively Lu et al. (2022). Details of the perturbative method will be provided in Sec. IV.1. In this work, we consider the following nonperturbative Hamiltonian:
@@ II.2 Nonperturbative interaction
where $s$ and $i$ are the spin and isospin indexes, respectively. The parameter $s_{\mathrm{L}}$ controls the range of the local part of the interaction. The smeared creation operator is defined as
@@
with $\tau_{I}$ the $I$th Pauli matrix .
The Hamiltonian in Eq. (13) consists of two parts, written separately for clarity. The first line contains terms that respect Wigner’s SU(4) symmetry Wigner (1937), an approximate symmetry of low-energy nuclear interactions. The inclusion of this symmetry helps suppress sign problems in lattice Monte Carlo calculations Chen et al. (2004). Recently, the SU(4)-symmetric Hamiltonian has been applied to the study of the properties of light and medium-mass nuclei as well as the equation of state of symmetric nuclear matter and neutron matter Lu et al. (2019), the intrinsic clustering geometry of 12C’s low-lying states Shen et al. (2023), the thermodynamics of nuclear systems Lu et al. (2020), the clustering effects in hot, dilute nuclear matter Ren et al. (2024), and the $\alpha$-particle monopole transition form factor Meißner et al. (2024). The second line of Eq. (13) introduces an isospin-dependent contact term and a OPE term, both of which explicitly break the SU(4) symmetry. These additional terms are introduced to improve the overall description of $p$-shell nuclei Wiringa and Pieper (2002) , such as 6He and 6Li studied in this work. To suppress the sign problem caused by the tensor force, we adopt a lower cutoff of $\Lambda_{\pi}^{\prime}=180$ MeV for the OPE term in Eq. (13), compared to the higher cutoff of $\Lambda_{\pi}=300$ MeV used in the full Hamiltonian in Eq. (1).
@@ III Axial Current
Figure 1: Diagrams illustrating the contributions to the axial current up to N3LO used in this work. Nucleons, pions and the external field are denoted by the solid , dashed, and wavy lines respectively. Panel (a) denotes the LO current while panel (b) is the relativistic correction at N2LO. Panels (c) and (d) are the OPE current and the contact current at N3LO, respectively.
@@
Since the momentum $\bm{k}$ carried by the current is very small in $\beta$-decay processes, we work in the $\bm{k}=0$ limit to simplify calculations. Under this limit, there are in total four terms that contribute to the axial current up to N3LO: the one-body current $\bm{A}_{\mathrm{LO}}$ at LO , the relativistic correction term $\bm{A}_{\mathrm{N}^{2}\mathrm{LO}}$ at N2LO, the OPE term $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$, and the contact term $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$ at N3LO. Figure 1 provides a diagrammatic illustration of them, and their expressions in the momentum space can be found in Refs. Baroni et al. (2016b); Park et al. (2003). The OPE current involves two nucleon-pion coupling constants, $c_{3}$ and $c_{4}$, for which we take the values $c_{3}=-5.61$ GeV-1 and $c_{4}=4.26$ GeV-1 Hoferichter et al. (2015). The contact current is related to the three-body force coefficient $c_{D}$ as follows Park et al. (2003); Baroni et al. (2018):
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})=-\frac{c_{D}}{8f_{\pi}^{2}\Lambda_{\chi}}\sum_{i<j}^{A}(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j}),$ (18)
where we used the shorthand notation $(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}=(\bm{\tau}_{i}\times\bm{\tau}_{j})^{x}+i(\bm{\tau}_{i}\times\bm{\tau}_{j})^{y}$ for the two-body isospin-raising operator.
The realization of the axial current on the lattice is straightforward. Given the momentum-space expression of the current, the finite lattice spacing and box size provides natural ultraviolet and infrared cutoffs, respectively, which are used to discretize the momenta and constrain them in the first Brillouin zone. The corresponding operator in the configuration space is then constructed through fast Fourier transformation.
As a comparison, in the following we briefly discuss the treatment of two-body currents for calculations in the continuum Park et al. (2003); Gazit et al. (2009); Baroni et al. (2018); King et al. (2020); Gnech et al. (2021). These calculations rely on analytical configuration-space expressions of the OPE and contact currents, which look like
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$ $=$ $\frac{g_{A}}{2f_{\pi}^{2}}\left(\frac{1}{3}c_{3}+\frac{2}{3}c_{4}+\frac{1}{6M_{N}}\right)\sum_{i<j}^{A}\delta^{(3)}(\bm{r}_{ij})$
$\times(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j})+\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$ $=$ $-\frac{c_{D}}{8f_{\pi}^{2}\Lambda_{\chi}}\sum_{i<j}^{A}\delta^{(3)}(\bm{r}_{ij})(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j}).$
The above equations can be easily derived by performing Fourier transformation on momentum-space expressions of the current. Note that $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$ consists of two parts: the first part proportional to $\delta^{(3)}(\bm{r}_{ij})=\delta^{(3)}(\bm{r}_{i}-\bm{r}_{j})$ absorbs the short-range singularity, while the second part with a tilde $\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) collects remaining long-range contributions, whose explicit form can be found in Refs. Pastore et al. (2018); King et al. (2020); Gnech et al. (2021). Because of the Fermi-Dirac statistics Park et al. (2003), the first part has the same spin-isopsin structure as the contact current $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$. Therefore, some works absorb the former into the latter and call the remaining term $\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) as the OPE current Park et al. (2003); Gazit et al. (2009); Baroni et al. (2018); King et al. (2020); Gnech et al. (2021). There is also literature which defines the full operator $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) as the OPE current Baroni et al. (2016b). We mention that these two definitions are equivalent to each other, and would lead to consistent predictions on physical observables such as the half-life. For our lattice calculation, we find it more convenient to follow the second prescription, which makes it easier to transform the current between the momentum representation and configuration representation, numerically.